Let a and b be the roots of x²-14x+36=0 show that a^n+b^n is divisible by 2n
1
Expert's answer
2020-02-05T08:51:51-0500

Solution. Solve the quadratic equation


"x^2-14x+36=0"

"D=(-14)^2-4\times36=52"

The roots of the equation is equal to


"x_1=\frac{14+\sqrt{52}}{2}=7+\sqrt{13}"

"x_2=\frac{14-\sqrt{52}}{2}=7-\sqrt{13}"

Consider the value


"(7+\sqrt{13})^k+(7-\sqrt{13})^k="

"=(7+\sqrt{13}+7-\sqrt{13})((7+\sqrt{13})^{k-1}+(7-\sqrt{13})^{k-1})-"


"-(7+\sqrt{13})(7-\sqrt{13})^{k-1}-(7-\sqrt{13})(7+\sqrt{13})^{k-1}="

"=14((7+\sqrt{13})^{k-1}+(7-\sqrt{13})^{k-1})-36((7+\sqrt{13})^{k-2}+(7-\sqrt{13})^{k-2})"

For the sum we got the recurrence equation


"S_k=14S_{k-1}-36S_{k-2}"

Solve the recurrence equation.


"S_1=7+\sqrt{13}+7-\sqrt{13}=14"

"S_2=2\times49+14\sqrt{13}-14\sqrt{13}+2\times13=124"

S1 is divisible by 2; S2 is divisible by 4.

Find S3


"S_3=14\times S_2-36\times S_1=14\times124-36\times14=1232"

1232 is not divisible by 6.